OpenStudy (anonymous):

7(x + 5) = 14, help?

OpenStudy (anonymous):

distribute out the 7 That means multiply 7 to x and 7 to 5 What do get then?

OpenStudy (anonymous):

I would get 7x+35=14?

OpenStudy (anonymous):

Yes you are correct. Then I would subtract 35 from both sides. Remember we want to get x by itself. What do we get then?

OpenStudy (anonymous):



OpenStudy (anonymous):

Did you already solve the question you posted?

OpenStudy (anonymous):

What do I do next? Do i subtract 14 on both sides?

OpenStudy (anonymous):

no you have to subtract 35 from both sides

OpenStudy (anonymous):



OpenStudy (anonymous):

Like that ^^^^

OpenStudy (anonymous):

7x+35=14 You got the equation wrong first off.

OpenStudy (anonymous):

7x+35-35=14 -35

OpenStudy (anonymous):

And then yuo cross out the 35s because they would equal o, right? And then you would get 7x=21 and then you would divide 7 on both sides and you would end up with 3, correct?

OpenStudy (anonymous):

7x=-21 So it will be -3 actually

OpenStudy (anonymous):

Good work though

OpenStudy (anonymous):

Oh ok, Thank you! =)